OpenStudy (anonymous):

Anyone good with Bernoulli Trials???? Assume that the coin is weighted so that a tail is 3 times as likely as a head. The coin is flipped 9 times. What is the probability that both heads and tails occur?___________________?

OpenStudy (anonymous):

\[1-\left(\left(\frac{3}{4}\right)^9+\left(\frac{1}{4}\right)^9\right)=\frac{60615}{65536} \]

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

I have tons of these